# DynamicGSG: Dynamic 3D Gaussian Scene Graphs for Environment Adaptation

### v1

@@ Abstract
In real-world scenarios, the environment changes caused by agents or human activities make it extremely challenging for robots to perform various long-term tasks. To effectively understand and adapt to dynamic environments, the perception system of a robot needs to extract instance-level semantic information, reconstruct the environment in a fine-grained manner, and update its environment representation in memory according to environment changes. To address these challenges, We propose DynamicGSG, a dynamic, high-fidelity, open-vocabulary scene graph generation system leveraging Gaussian splatting. Our system comprises three key components: (1) constructing hierarchical scene graphs using advanced vision foundation models to represent the spatial and semantic relationships of objects in the environment, (2) designing a joint feature loss to optimize the Gaussian map for incremental high-fidelity reconstruction, and (3) updating the Gaussian map and scene graph according to real environment changes for long-term environment adaptation. Experiments and ablation studies demonstrate the performance and efficacy of the proposed method in terms of semantic segmentation, language-guided object retrieval, and reconstruction quality. Furthermore, we have validated the dynamic updating capabilities of our system in real laboratory environments. The source code will be released at: https://github.com/GeLuzhou/DynamicGSG.
@@ I Introduction
Future intelligent robots are supposed to execute diverse long-term complex tasks in dynamic environments based on intricate natural language instructions from humans. To achieve this, agents must possess dynamic environment perception and comprehension capabilities. Prior studies [1, 2],[3, 4] construct static open-vocabulary scene graphs from sensor data to capture the topological structures of environments at specific moments, which enhances robots’ understanding of complex instructions and facilitates task completion. However, static scene graphs are of limited use in real-world scenarios, as robot workspaces typically change due to human activities or other agents’ operations. Additionally, The inherent latency between the scene graph stored in memory (which guides task planning and execution) and the actual environment state significantly hinders successful task completion.
Recent advancements in 3D Gaussian splatting[5],[6],[7],[8],[9] have attracted significant attention within the robotics community. These developments have applications in various domains, including high-quality reconstruction[5],[6], SLAM[10, 11], robotic manipulation[12], semantic embedding[13, 14], and 3D environment understanding[15, 16]. The explicit point-based representation of 3D Gaussians effectively facilitates the integration of semantic information from advanced vision foundation models, enabling the construction of topological scene representations. Based on fast differentiable rendering capability, we find 3D Gaussians are particularly suitable for rapid updates of reconstructed scenes. Moreover, the high-fidelity environment reconstruction provided by 3D Gaussians naturally supports the development of dynamic, high-quality scene graph construction systems.
@@
Most previous works on scene graph construction primarily rely on point clouds[1, 2],[4] or meshes [17, 18]. These methods often struggle to promptly respond to dynamic environment changes and fail to capture fine-grained details of 3D scenes due to the inherent limitations of traditional representation techniques.
@@ II Related Work
3D scene graph[22, 18] offers a hierarchical graph-structured representation of the environment with nodes representing spatial concepts at multiple levels of abstraction (e.g., objects, places, rooms, buildings.) and edges preserving the spatial and semantic relationships between nodes. Taking advantage of the generalization abilities of vision foundation models[19, 20, 23, 24] and cross-modal grounding capabilities of vision-language models[21, 25] enables the construction of 3D scene graphs at the open-vocabulary level. Recent methods [1, 2],[4, 3, 26] construct 3D scene graphs embedded with VLM semantic features, facilitating object retrieval, robot manipulation and navigation.
@@ II-B Gaussian-based Open-Vocabulary Scene Understanding
Recent progress in Gaussian Splatting [5, 6],[7, 8],[9] demonstrates outstanding performance in photo-realistic reconstruction with promising efficiency. The gradient-based optimization has also been applied in an online setting to incrementally construct the Gaussian map[10, 11],[28, 29] through differentiable rendering.
@@
To address these deficiencies, we incrementally construct the semantic Gaussian map using posed RGB-D sequences from public datasets or live cameras running Visual-Inertial Odometry (VIO) [31, 32],[33, 34] frameworks, enabling it to achieve high-fidelity reconstruction and integrate rich semantic information to build a topological scene graph. Leveraging the rapid training and differentiable rendering capabilities of 3D Gaussians, our system can dynamically update both the Gaussian maps and scene graphs to adapt to changes in the real-world environments.
@@ III-A Gaussian Preliminaries
We densely represent the scene using isotropic Gaussian, which is an explicit representation parameterized by RGB color $\mathbf{c}$, center position $\bm{\mu}\in\mathbb{R}^{3}$, radius $\mathbf{r}\in\mathbb{R}^{+}$, and opacity $\mathbf{o}\in[0,1]$. The influence of each Gaussian on 3D space point $\mathbf{x}\in\mathbb{R}^{3}$ is defined as:
$f(\mathbf{x})=\mathbf{o}\cdot\exp\left(-\frac{\|\mathbf{x}-\bm{\mu}\|^{2}}{2\mathbf{r}^{2}}\right)$ (1)
@@ III-B Gaussian Objects Association
For all detections $D_{t}=\{\textbf{d}_{t,i}\}_{i=1}^{M}$ in $I_{t}$, we associate them with the map objects $O_{t-1}$ through joint geometric and semantic similarity matching. As shown in Fig. 3, we first render all Gaussians into the current camera view at instance-level to obtain object masks $\{\textbf{m}_{t,o_{j}}\}_{j=1}^{J}$. Objects $O_{t-1}$ whose projected masks contain more than $\delta_{\text{pix}}$ pixels in the image plane are considered visible in the current frame. The geometric similarity is then formulated by the Intersection over Union (IoU) between all 2D detection masks $\{\textbf{m}_{t,i}\}_{i=1}^{M}$ and projected masks of visible 3D objects $\{\textbf{m}_{t,o_{j}}\}_{j=1}^{J}$:
@@ IV-A Experiment Setups
To generate the semantic segmentation, we first calculate the CLIP text description vector for class-specific prompt formatted as “an image of {class label}" corresponding to each class in the Replica dataset. For each scene, we compute the cosine similarity between the semantic feature of each object within the scene graph and the text description vector of each class. Each object’s points or Gaussians are allocated to the class with the highest similarity score. Finally, the point clouds or Gaussians generated by all methods are transformed to the same coordinate as the ground-truth semantic point clouds. Quantitative evaluation is performed through standardized metrics, including mAcc, mIoU, and frequency-weighted mIoU.
As shown in Tab. I, our method performs better than all baseline methods on mAcc and mIoU while achieving comparable performance to HOV-SG on F-mIoU. And the joint feature loss also effectively regularizes the Gaussian instance grouping to improve the mIOU and F-mIOU. The qualitative results in Fig. 4 further demonstrate that joint feature loss significantly enhances the regularization of intra-instance Gaussian grouping.
@@ IV-C Language-guided Object Retrieval
To validate whether the multi-layer scene graphs constructed by DynamicGSG can effectively capture spatial and semantic object relationships, we conduct a language-guided object retrieval experiment employing diverse query types across three semantic complexity levels (Descriptive: E.g., “These are some books and they are on the table."; Affordance: E.g., “Something I can open with my keys."; Negation: E.g., “Something to sit on other than a chair.") provided by ConceptGraphs[1] on the Replica dataset[36]. Following ConceptGraphs, we select 20 Descriptive, 5 Affordance and Negation queries for each scene, ensuring each query corresponds to at least one ground-truth object.
@@
Analysis of top-1, top-2 and top-3 recall across diverse query types in Tab. IV-A indicates that hierarchical matching substantially enhances accuracy for descriptive queries involving inter-object relationships, exhibiting a significant improvement over ConceptGraphs[1]. This performance improvement is qualitatively illustrated in Fig. 5, where our hierarchical matching approach successfully localizes target objects that baseline fails to identify. Futhermore, LLMs demonstrate superior instruction comprehension for Affordance and Negation queries, which more closely approximate natural human language. DynamicGSG also facilitates earlier object localization within these query categories.
@@ IV-E Real-world Dynamic Update
As detailed in Tab. V, DynamicGSG, while leveraging the method detailed in Sec. III-E to perform initial pose refinement and incrementally construct Gaussian scene graphs, successfully detects environment changes and executes corresponding dynamic updates, ensuring temporal consistency between the scene graphs and the real-world environments. Some visualization results of the experiment, presented in Figures 1 and 6, demonstrate DynamicGSG detects three types of environment changes and effectively performs instance-level local updates to construct dynamic, high-fidelity Gaussian scene graphs.

### v2

@@ Abstract
In real-world scenarios, environment changes caused by human or agent activities make it extremely challenging for robots to perform various long-term tasks. Recent works typically struggle to effectively understand and adapt to dynamic environments due to the inability to update their environment representations in memory according to environment changes and lack of fine-grained reconstruction of the environments. To address these challenges, we propose DynamicGSG, a dynamic, high-fidelity, open-vocabulary scene graph construction system leveraging Gaussian splatting. DynamicGSG builds hierarchical scene graphs using advanced vision language models to represent the spatial and semantic relationships between objects in the environments, utilizes a joint feature loss we designed to supervise Gaussian instance grouping while optimizing the Gaussian maps, and locally updates the Gaussian scene graphs according to real environment changes for long-term environment adaptation. Experiments and ablation studies demonstrate the performance and efficacy of our proposed method in terms of semantic segmentation, language-guided object retrieval, and reconstruction quality. Furthermore, we validate the dynamic updating capabilities of our system in real laboratory environments. The source code and supplementary experimental materials will be released at: https://github.com/GeLuzhou/Dynamic-GSG.
@@ I Introduction
Future intelligent robots are supposed to execute diverse long-term complex tasks in dynamic environments based on intricate natural language instructions from humans. To achieve this, agents must possess dynamic environment perception and comprehension capabilities. Prior studies [1, 2, 3, 4] construct static open-vocabulary scene graphs from sensor data to capture the topological structures of environments at specific moments, which enhances robots’ understanding of complex instructions and facilitates task completion. However, static scene graphs are of limited use in real-world scenarios, as robot workspaces typically change due to human activities or other agents’ operations. Additionally, the inherent latency between the scene graph stored in memory (which guides task planning and execution) and the actual environment state significantly hinders successful task completion.
Recent advancements in 3D Gaussian splatting [5, 6, 7, 8] have attracted significant attention within the robotics community. These developments have applications in various domains, including high-quality reconstruction [5, 6, 9, 10] robotic manipulation [11], semantic embedding [12, 13], and 3D environment understanding [14, 15]. The explicit point-based representation of 3D Gaussians effectively facilitates semantic information integration from advanced vision language models, enabling the construction of topological scene representations. Based on fast differentiable rendering and explicit representation, we find 3D Gaussians are particularly suitable for locally rapid updates of reconstructed scenes. Moreover, the high-fidelity environment reconstruction provided by 3D Gaussians naturally supports the development of dynamic, high-quality scene graph construction systems.
@@
Most previous works on scene graph construction primarily rely on point clouds [1, 2, 4, 16, 17] These methods often struggle to promptly respond to dynamic environment changes and fail to capture fine-grained details of 3D scenes due to the inherent limitations of traditional representation techniques.
@@ II Related Work
3D scene graph [21, 17] offers a hierarchical graph-structured representation of the environment with nodes representing spatial concepts at multiple levels of abstraction (e.g., objects, places, rooms, buildings.) and edges preserving the spatial and semantic relationships between nodes. Taking advantage of the generalization abilities of vision foundation models [18, 19, 22, 23] and cross-modal grounding capabilities of vision-language models [20, 24] enables the construction of 3D scene graphs at the open-vocabulary level. Recent methods [1, 2, 4, 3, 25] construct 3D scene graphs embedded with VLM semantic features, facilitating object retrieval, robot manipulation and navigation.
@@ II-B Gaussian-based Open-Vocabulary Scene Understanding
Recent progress in Gaussian Splatting [5, 6, 7, 27, 8] demonstrates outstanding performance in photo-realistic reconstruction with promising efficiency. The gradient-based optimization has also been applied in an online setting to incrementally construct the Gaussian map [9, 10, 28, 29] through differentiable rendering.
@@
To address these deficiencies, we incrementally construct the semantic Gaussian map using posed RGB-D sequences from public datasets or live cameras running Visual-Inertial Odometry (VIO) [31, 32, 33, 34] frameworks, enabling it to achieve high-fidelity reconstruction and integrate rich semantic information to build a topological scene graph. Leveraging the rapid training and differentiable rendering capabilities of 3D Gaussians, our system can dynamically update both the Gaussian maps and scene graphs to adapt to changes in the real-world environments.
@@ III-A Gaussian Preliminaries
We densely represent the scene using isotropic Gaussian, which is an explicit representation parameterized by RGB color $\mathbf{c}$, center position $\boldsymbol{\mu}\in\mathbb{R}^{3}$, radius $\mathbf{r}\in\mathbb{R}^{+}$, and opacity $\mathbf{o}\in[0,1]$. The influence of each Gaussian on 3D space point $\mathbf{x}\in\mathbb{R}^{3}$ is defined as:
$f(\mathbf{x})=\mathbf{o}\cdot\exp\left(\frac{\|\mathbf{x}-\boldsymbol{\mu}\|^{2}}{2\mathbf{r}^{2}}\right)$ (1)
@@ III-B Gaussian Objects Association
For all detections $D_{t}=\{\textbf{d}_{t,i}\}_{i=1}^{M}$ in $I_{t}$, we associate them with the map objects $O_{t-1}$ through joint geometric and semantic similarity matching. As shown in we first render all Gaussians into the current camera view at instance-level to obtain object masks $\{\textbf{m}_{t,o_{j}}\}_{j=1}^{J}$. Objects $O_{t-1}$ whose projected masks contain more than $\delta_{\text{pix}}$ pixels in the image plane are considered visible in the current frame. The geometric similarity is then formulated by the Intersection over Union (IoU) between all 2D detection masks $\{\textbf{m}_{t,i}\}_{i=1}^{M}$ and projected masks of visible 3D objects $\{\textbf{m}_{t,o_{j}}\}_{j=1}^{J}$:
@@ IV-A Experiment Setups
To generate the semantic segmentation, we first calculate the CLIP text description vector for class-specific prompt formatted as “an image of {class label}” corresponding to each class in the Replica dataset. For each scene, we compute the cosine similarity between the semantic feature of each object within the scene graph and the text description vector of each class. Each object’s points or Gaussians are allocated to the class with the highest similarity score. Finally, the point clouds or Gaussians generated by all methods are transformed to the same coordinate as the ground-truth semantic point clouds. Quantitative evaluation is performed through standardized metrics, including mAcc, mIoU, and frequency-weighted mIoU.
As shown in Tab. I, our method performs better than all baselines on mAcc and mIoU while achieving comparable performance to HOV-SG on F-mIoU. Object association in [1, 2] relies on the overlap ratio between point clouds which suffers from a critical limitation: the potential association of small objects into nearby large objects due to high overlap ratios of 3D point clouds. Our method employs 3D-2D object association which effectively prevents spurious merging, as the small masks and large objects do not exhibit abnormal geometric similarity, ultimately yielding a significant enhancement in mAcc. And the joint feature loss also effectively regularizes the Gaussian instance grouping to improve the mIOU and F-mIOU. The qualitative results in Fig. 4 further demonstrate that joint feature loss significantly enhances the regularization of intra-instance Gaussian grouping.
@@ IV-C Language-guided Object Retrieval
To validate whether the multi-layer scene graphs constructed by DynamicGSG can effectively capture spatial and semantic object relationships, we conduct a language-guided object retrieval experiment employing diverse query types across three semantic complexity levels (Descriptive: E.g., “These are some books and they are on the table.”; Affordance: E.g., “Something I can open with my keys.”; Negation: E.g., “Something to sit on other than a chair.”) provided by ConceptGraphs [1] on the Replica dataset [36]. Following ConceptGraphs, we select 20 Descriptive, 5 Affordance and Negation queries for each scene, ensuring each query corresponds to at least one ground-truth object.
@@
Analysis of top-1, top-2 and top-3 recall across diverse query types in Tab. IV-A indicates that hierarchical matching substantially enhances accuracy for descriptive queries involving inter-object relationships, exhibiting a significant improvement over ConceptGraphs [1]. This performance improvement is qualitatively illustrated in Fig. 5, where our hierarchical matching approach successfully localizes target objects that baseline fails to identify. Futhermore, LLMs demonstrate superior instruction comprehension for Affordance and Negation queries, which more closely approximate natural human language. DynamicGSG also facilitates earlier object localization within these queries.
@@ IV-E Real-world Dynamic Update
As detailed in Tab. V, DynamicGSG, while leveraging the method detailed in Sec. III-E to perform initial pose refinement and incrementally construct Gaussian scene graphs, successfully detects environment changes and executes corresponding dynamic updates, ensuring temporal consistency between the scene graphs and the real-world environments. Some visualization results of the experiment, presented in Fig. 1 and 6, demonstrate DynamicGSG detects three types of environment changes and effectively performs instance-level local updates to construct dynamic, high-fidelity Gaussian scene graphs.